OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). ( -2i is a zero of f(x) = x^4 - 5x^2 - 36 These are the answer choices: (a)2i, 6i, -6i (b)2i, 6, -6 (c)2i, 3, -3 (d)2i, 3i, -3i

OpenStudy (danjs):

if -2i is a zero, complex numbers come in pairs so +2i will also be a zero

OpenStudy (danjs):

(a+2i)(a-2i)

OpenStudy (danjs):

so those are 2 of the terms for sure

OpenStudy (danjs):

there are 2 more, since it is a 4th order polynomial

OpenStudy (anonymous):

\[ x^4 - 5x^2 - 36=(x+2i)(x-2i)(something)=(x^2+4)(something)\] find the "something by factoring

OpenStudy (anonymous):

or you can start by factoring \[ x^4 - 5x^2 - 36\]

OpenStudy (anonymous):

(x^2+4) (x-3) (x+3)

OpenStudy (danjs):

9 * 4 = 36 and 4 - 9 = -5

OpenStudy (anonymous):

?

OpenStudy (danjs):

Yes!

OpenStudy (anonymous):

awesome :D

OpenStudy (danjs):

so each of those (x - 2i)(x +2i)(x-3) (x+3) = 0 If one of those terms is 0, then the equation is true.

OpenStudy (anonymous):

what terms?

OpenStudy (danjs):

When you factor f(x) = x^4 - 5x^2 - 36 , you get (x - 2i)(x +2i)(x-3) (x+3) = 0 for that to be true (x-2i) = 0 or (x+2i) = 0 or (x-3) = 0 or (x+3) = 0

OpenStudy (danjs):

so the 4 possible answers or zeros of f(x) are.....

OpenStudy (anonymous):

Idk how could i reach the answer?

OpenStudy (anonymous):

@DanJS